an infiniti G35 travels at an average speed of 50 miles per hour. A Mercury Mountaineer travels at an average speed of 40 miles per hour. In the time it takes the Mountaineer to travel a certain distance d, the Infiniti travels 100 miles farther. Find the distance that each car travels and the time of the trip.
thank You for any and all help you can give me. Can't figure out how to set this up.

Answer by scott8148(6628)

there are two equations and two unknowns (distance and time)
__ solving by substitution is probably the best technique
__ d=rt

G35 __ d+100=50t

Mountaineer __ d=40t

substituting __ (40t)+100=50t __ 100=10t __ 10=t

substituting __ d=40(10) __ d=400